OpenStudy (anonymous):

Determine the discriminant of x2 = 4.

OpenStudy (aravindg):

For a quadratic ax^2+bx+c=0 Discriminant=b^2-4ac

OpenStudy (anonymous):

Determine the discriminant of x2 = 4. possible answers: 17 16 -15 -10

OpenStudy (aravindg):

For a quadratic ax^2+bx+c=0 Discriminant=b^2-4ac Use that

OpenStudy (anonymous):

thank you i got my answer

OpenStudy (aravindg):

Yw :) ...Click best response if someone's answer helped you.

OpenStudy (anonymous):

idk this one though: Determine the discriminant of 2x2 − 3x + 1 = 0. possible answers: 15 -17 1 4